OpenStudy (anonymous):

Given the functions h:x: 4x + m and h^-1:x : 2kx + 5/8 , where m and k are constants, find the value of m and of k.

OpenStudy (anonymous):

@crazysingh can you help me?

OpenStudy (anonymous):

wait i am trying .

OpenStudy (anonymous):

is \[m = \frac{ -5 }{ 2 } ; k = \frac{ 1 }{ 8 }\] am i correct ?

OpenStudy (anonymous):

whoa, correct!

OpenStudy (anonymous):

ok i will explain.

OpenStudy (anonymous):

first find inverse of h(x) by substituting x by y and y by and solve for y = f(x). f(x) will be inverse of the given function h(x). \[y= 4x+m\] so \[x = \frac{ y-m }{ 4 } = \frac{ 2y-2m }{ 8 }\] now substitute x by y and y by x to get the inverse. \[y = \frac{ 2x-2m }{ 8 }\] .. [I] above equation represents inverse of the function. compare it with given inverse function. which is \[h ^{-1}(x) = 2kx + \frac{ 5 }{ 8 } = \frac{ 16kx+5 }{8 }\] .. [II] compare [I] and [II] to get values of k and m.

OpenStudy (anonymous):

let me try it first.

OpenStudy (anonymous):

why must the equation times with 2